OpenStudy (anonymous):

What is the simplified form of -3x-6x+x?

OpenStudy (krabsen):

-8x

OpenStudy (lgbasallote):

what is -3 - 6 + 1?

OpenStudy (krabsen):

-3-6=-9 -9+1=-8

OpenStudy (anonymous):

Ohh i get. I'm not good with simplifying equations,lol

OpenStudy (anonymous):

Would you guys mind helping me on a couple more?

OpenStudy (krabsen):

that is ok but im not going to make all of them for you :)

OpenStudy (krabsen):

you make them, and i will tell what you do wrong, if you do anything wrong

OpenStudy (anonymous):

Ok, sounds good to me.

OpenStudy (anonymous):

Ok what about 2x+3y-4x Is it 6x+3y?

OpenStudy (krabsen):

no the last 4x are negativ

OpenStudy (krabsen):

so that is 2x-4x

OpenStudy (anonymous):

Would it be 2x?

OpenStudy (anonymous):

-2x?

OpenStudy (krabsen):

yes -2x

OpenStudy (krabsen):

so the anwser is -2x+3y

OpenStudy (anonymous):

What about (2x+3y-5)+(-3y-6

OpenStudy (anonymous):

I knew it thank you.

OpenStudy (krabsen):

first you need to remove ( ), im danish so not all i can spell in english

OpenStudy (anonymous):

But they are two seperate equations, wouldn;t you have to distribute?

OpenStudy (krabsen):

there are not two equations here

OpenStudy (krabsen):

you just need to left the ( ), and then add x's and y's

OpenStudy (anonymous):

So would it be 2x-1y-1?

OpenStudy (krabsen):

you got 3y and -3y that is not 1y

OpenStudy (anonymous):

I'm confused.

OpenStudy (krabsen):

(2x+3y-5)+(-3y-6) 2x+3y-5-3y-6 2x-11

OpenStudy (krabsen):

2x=2x 3y-3y=0 -5-6=-11

OpenStudy (anonymous):

Ohhhh, ok. =)

OpenStudy (anonymous):

4x+3y)-(2x-5y)?

OpenStudy (krabsen):

it is the same, but this time the 2ed ( ) is negativ

OpenStudy (anonymous):

I got 2x+8y

OpenStudy (krabsen):

yes

OpenStudy (anonymous):

What about 3xy+4xy-6xz

OpenStudy (krabsen):

here you add xy's

OpenStudy (anonymous):

The answer is 7xy-6xz

OpenStudy (krabsen):

yes

OpenStudy (anonymous):

What about 7xy+2xz+4xz-12xyz+2zxt

OpenStudy (anonymous):

**7xy+2xz+4xz-12xyz+2zx

OpenStudy (krabsen):

here you find all the ones with the same letters after the number and add them

OpenStudy (anonymous):

Is the answer 7xy+8xz-12xyz?

OpenStudy (krabsen):

yes

OpenStudy (anonymous):

What about -3(2x-5)-4(x+7)

OpenStudy (krabsen):

here you need to mulitply to lift ( )

OpenStudy (anonymous):

would it be -10x-13?

OpenStudy (krabsen):

yes

OpenStudy (anonymous):

Thank you, that is all the problems I have, you were a great help, I hope you can continue hlepingme and others the same =)

OpenStudy (krabsen):

np happy to help :)

OpenStudy (anonymous):

=)